OpenStudy (anonymous):

@jim_thompson5910 how many committees of four can be chosen from twelve students? how many of these will include a given student? how many will exclude a given student?

jimthompson5910 (jim_thompson5910):

how many committees of four can be chosen from twelve students? use n C r = (n!)/(r!*(n-r)!) where n = 12 and r = 4

OpenStudy (anonymous):

yes i get the first question. but what about hte second and third?

jimthompson5910 (jim_thompson5910):

well if a student must be on the committee, then you have 4-1 = 3 slots left and you have 12-1 = 11 people to fill those 3 slots

jimthompson5910 (jim_thompson5910):

does that help?

OpenStudy (anonymous):

how did u know that?

jimthompson5910 (jim_thompson5910):

you have 4 slots to start with

jimthompson5910 (jim_thompson5910):

a certain student (say student X) must be on this committee

jimthompson5910 (jim_thompson5910):

So you go from _ _ _ _ to X _ _ _

jimthompson5910 (jim_thompson5910):

that's why you have 3 slots left and 11 people left to fill them

OpenStudy (anonymous):

ok so that's the second question then what's third quesiton

jimthompson5910 (jim_thompson5910):

parts b and c are the complete opposite

jimthompson5910 (jim_thompson5910):

if you know how many ways to order a committee with student X on the committee, then you can find out how many ways to have student X off of the committee

OpenStudy (kropot72):

If one student is excluded there are only 11 to choose from. 11C4

OpenStudy (anonymous):

the answer to these questions were 495, 165, 330

jimthompson5910 (jim_thompson5910):

a) 12 C 4 = 495 b) 11 C 3 = 165 c) 11 C 4 = 330 (notice how 11 C 3 + 11 C 4 = 12 C 4) so that's how they got those answers

OpenStudy (anonymous):

i get the first two questions rihgt... but number three im still confused

jimthompson5910 (jim_thompson5910):

basically the numbers from parts b and c must add up to the total of 495

jimthompson5910 (jim_thompson5910):

so that's why 165 + 330 = 495

OpenStudy (kropot72):

Do you know how to calculate 11C4 ? This gives the answer to number 3 question.

OpenStudy (anonymous):

yes i do. but getting 11C4 im stil thinking about it

OpenStudy (anonymous):

like why not 12c4? like how u guys know its 11?

jimthompson5910 (jim_thompson5910):

because you can't pick the student X if you're keeping him/her off the committee

jimthompson5910 (jim_thompson5910):

you have 12-1 = 11 choices left but you still have 4 slots

OpenStudy (anonymous):

haha alrhgt thank you i get it now!

jimthompson5910 (jim_thompson5910):

ok that's great

OpenStudy (anonymous):

lol omg haha lol ofcourse its 11 cuz u kick him off tahnks dude!

jimthompson5910 (jim_thompson5910):

yw